# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

### v3

@@ Appendix F Compute Resources
Appendix G Configurations of Tasks
We provide the detailed configurations of the tasks in the experiments in this section.
@@ Appendix G Configurations of Tasks
  • AntSpeed: The ant robot is trained to move as fast as possible. The reward is given if the ant robot reaches a target speed. Maximum episode length is 1000 steps.
@@
  • AntVeryFar: The ant robot is trained to reach a target position very far from the starting point. The reward is given if the ant robot reaches the target position. Maximum episode length is 1000 steps.
  • WalkerKeep: The walker robot is trained to keep a target height. The robot is initialized by standing, and the reward is given if the robot maintains the target height. Maximum episode length is 1000 steps.

### v4

@@ A.6.3 Compute Resources
A.7 Configurations of Tasks
In this section, we provide the detailed configurations of the tasks in the experiments.
@@ A.7 Configurations of Tasks
  • RobotSlide: The robot arm is trained to slide an object to a target position. The target position is randomly generated on the table, and the reward is given if the object reaches the target position. Maximum episode length is 500 steps.
  • RobotPickPlace: The robot arm is trained to pick and place an object to a target position. The target position is randomly generated in the space, and the reward is given if the object reaches the target position. Maximum episode length is 500 steps.
  • Pitfall: The agent is tasked with collecting all the treasures in a jungle while avoiding the pitfalls. The reward is given if the agent collects one treasure, while if the agent falls into a pitfall, the episode ends. Maximum episode length is 2000 steps.
  • Frogger: The agent is trained to cross frogs on a river. The reward is given when each frog is crossed, and the episode ends if all frogs are crossed or fall into the river. Maximum episode length is 2000 steps.
  • MontezumaRevenge: The agent is trained to navigate through a series of rooms to collect keys and reach the final room. The reward is given if the agent successfully reaches one new room. Maximum episode length is 5000 steps.
  • Solaris: The agent controls a spaceship to blast enemies and explore new galaxies. The reward is given if the agent destroys one enemy spaceship and enters a new galaxy. Maximum episode length is 2000 steps.
  • Freeway: The agent is trained to guide the chicken across multiple lanes of heavy traffic. The reward is given if one chicken crosses one lane, while the episode ends if all chickens are crossed or hit by a car. Maximum episode length is 2000 steps.
@@
Furthermore, we provide the detailed dimensions of the states in our evaluated tasks in Table 14.
Table 14: The dimensions of the states in the evaluated tasks.
Domain (Tasks) Dimension
Ant robot (AntStand, AntFar) 105
Humanoid robot (HumanStand, HumanKeep) 348
RobotReach 20
RobotPush, RobotSlide and RobotPickPlace 35
Atari games (MontezumaRevenge, PitFall, Frooger, Solaris, Freeway) $84\times 84=7056$
MountainCar 2